40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 819     Entry time: Fri Feb 10 23:47:22 2012
Author: Tara, Frank 
Type: DailyProgress 
Category: BEAT 
Subject: beam splitter mount replaced 

we replaced the mount for the combining beam splitter in the beat setup as it caused a large, broadband peak in the spectrum around 1.4kHz. The new mount is one of the old, fixed turning mirror blocks they used in initial LIGO at LLO as far as i know. After replacing the mount the peak is entirely gone. I've used two springs instead of one to increase the pressure.  We could not determine the resonance frequency of the new mount. Tapping the mount excites only known mechanical resonances from the surrounding mirror mounts. Tara posted a plot for comparison before and after replacing that mount (see here). He also has prepared a nice plot combined with a drawing which mount corresponds to which resonance we see in the spectrum. We will use this to start reducing (or even eliminating) those resonances starting with the most dominant ones close to 1kHz

Attached a copy of the drawing.

SCAN1623_000.pdf

 

ELOG V3.1.3-